OpenStudy (anonymous):

N-1/2>3/4 the greater than equal to sign has a line under it my iPod wouldn't let me put that sooo

OpenStudy (anonymous):

N-1/2 is greater than or equal to 3/4 so N is greater than or equal to 1/4.

OpenStudy (anonymous):

18) q+0.5>-2 solve the inequality 20)5x<15 22) y/3<2 there is a line under the < 24) 25.5g>102 All these are inequality thank you for your help

OpenStudy (anonymous):

\[N-1/2 \ge 3/4 \]

OpenStudy (anonymous):

I won't do all your work for you, but you need to get X or what ever the variable is to one side to solve the inequality. For number 20 it would be 5x<15 Divide by 5 on both sides x<3

OpenStudy (anonymous):

Yeah well how can I possibly get q+0.5>-2 how do I get q to the other side

OpenStudy (anonymous):

q+0.5>-2 Subtract 0.5 from both sides q>-2.5

OpenStudy (anonymous):

So for numer 22) I divide the 2 into y/3

OpenStudy (anonymous):

You want Y on its own side, so you multiply the 3 into the 2 to get rid of the 3 under the Y because \[\frac{ y }{ 3 } X \frac{ 3 }{ 1 } = y \] The 3's cancel out leaving Y on its own side.

OpenStudy (anonymous):

Ohh so I was wrong